|x - 2| + 3 = 10
im so confused.
Ive tried over and over, please help!

Answer by Edwin McCravy(20086)

|x - 2| + 3 = 10
im so confused.
Ive tried over and over, please help!
|x - 2| + 3 = 10 

Subtract 3 from both sides:

|x - 2| + 3 = 10
        - 3  - 3
----------------
|x - 2|     =  7

To remove the absolute value bars make two equations:

x - 2 = 7          x - 2 = -7

Solve each:

    x = 9             x = -5

Those are the two solutions.

Edwin